### v1

# Fine-Grained Motion Compression and Selective Temporal Fusion for Neural B-Frame Video Coding

Abstract
With the remarkable progress in neural P-frame video coding, neural B-frame coding has recently emerged as a critical research direction. However, most existing neural B-frame codecs directly adopt P-frame coding tools without adequately addressing the unique challenges of B-frame compression, leading to suboptimal performance. To bridge this gap, we propose novel enhancements for motion compression and temporal fusion for neural B-frame coding. First, we design a fine-grained motion compression method. This method incorporates an interactive dual-branch motion auto-encoder with per-branch adaptive quantization steps, which enables fine-grained compression of bi-directional motion vectors while accommodating their asymmetric bitrate allocation and reconstruction quality requirements. Furthermore, this method involves an interactive motion entropy model that exploits correlations between bi-directional motion latent representations by interactively leveraging partitioned latent segments as directional priors. Second, we propose a selective temporal fusion method that predicts bi-directional fusion weights to achieve discriminative utilization of bi-directional multi-scale temporal contexts with varying qualities. Additionally, this method introduces a hyperprior-based implicit alignment mechanism for contextual entropy modeling. By treating the hyperprior as a surrogate for the contextual latent representation, this mechanism implicitly mitigates the misalignment in the fused bi-directional temporal priors. Extensive experiments demonstrate that our proposed codec outperforms state-of-the-art neural B-frame codecs and achieves comparable or even superior compression performance to the H.266/VVC reference software under random-access configurations.
Index Terms:
B-Frame Video Coding, Motion Compression, Neural Networks, Neural Video Coding, Temporal Fusion.I Introduction
The dramatic growth of video traffic has created unprecedented demands for efficient video coding technologies. To improve compression efficiency, a series of video coding standards have been developed over the past few decades, including H.264/AVC [1], H.265/HEVC [2], and H.266/VVC [3]. While these standards have achieved remarkable success, further advancements in compression performance remain challenging.
Recently, neural video coding has gained significant attention [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. Existing schemes primarily focused on two scenarios: low-delay coding and random-access coding. In the low-delay scenario, the current frame (P-frame) is constrained to reference only previously encoded frames, ensuring minimal latency. In contrast, in the random-access scenario, the current frame (B-frame) is allowed for bi-directional referencing, trading latency for higher compression efficiency. Current research efforts concentrate on two aspects: (1) how to obtain compact yet accurate motion representation and (2) how to effectively utilize predicted temporal contexts. For the first challenge, for example, the methods such as multiframe-based motion vector prediction [10], resolution-adaptive motion vector compression [8], and coarse-to-fine motion vector compression [5] have effectively improved the accuracy of motion-compensated prediction and the efficiency of motion compression. For the second challenge, innovations such as condition coding [27] have shown advantages over traditional residual coding by better exploiting temporal correlations. However, most existing neural B-frame coding solutions [33, 32, 34] simply adapt these P-frame coding tools without fully considering the unique characteristics of bi-directional prediction, which limits their performance potential.
B-frame coding exhibits two distinctive characteristics that pose unique challenges. First, it requires estimating bi-directional motion vectors from bi-directional reference frames, leading to higher motion coding costs compared to P-frame coding. Current schemes typically process bi-directional motion vectors by either: (1) concatenating them for joint compression using a single motion auto-encoder [33, 32], or (2) employing a parameter-shared motion auto-encoder for separate compression [34]. However, these coarse-grained motion vector compression methods overlook two fundamental aspects of bi-directional motion characteristics: (a) the temporal correlation asymmetry between forward and backward references leads to distinct bitrate allocation needs and reconstruction quality requirements for bi-directional motion vectors, and (b) the inherent geometric consistency between forward and backward motion vectors can be exploited through joint motion modeling to improve compression efficiency. Second, the quality of predicted bi-directional temporal contexts and temporal priors varies substantially due to the differences in temporal reference correlation and motion vector precision. Existing schemes [33, 32, 34] fused these contexts and priors uniformly through the contextual encoder-decoder and contextual entropy model without quality discrimination, which may propagate prediction errors and ultimately degrade rate-distortion performance. These characteristics suggest that dedicated algorithmic designs for B-frame coding may be necessary to achieve further compression efficiency improvement.
In this paper, to reduce the bi-directional motion coding costs, we propose a fine-grained motion compression method for neural B-frame video coding. Specifically, we propose an interactive dual-branch motion auto-encoder to separately compress and reconstruct forward and backward motion vectors. This motion auto-encoder uses a cross-branch motion information interaction mechanism and per-branch adaptive quantization steps to facilitate bi-directional motion information exchange and fine-grained motion rate-distortion control. Furthermore, we propose an interactive motion entropy model that separately estimates the bitrate for each directional motion latent representation. This entropy model interactively leverages partitioned segments of bi-directional motion latent representations as direction priors, thereby effectively capturing fine-grained bi-directional motion dependencies and enhancing compression efficiency.
To enable discriminative utilization of bi-directional temporal contexts and temporal priors, we propose a selective temporal fusion method for neural B-frame video coding. Specifically, during contextual encoding and decoding, our contextual encoder-decoder predicts bi-directional fusion weights, which serve as discriminative cues to guide the selective fusion of multi-scale bi-directional temporal contexts. Beyond the contextual encoder-decoder, we further introduce selective temporal prior fusion in the entropy modeling stage. By treating the hyperprior as a surrogate for the contextual latent representation, we implicitly align the fused bi-directional temporal priors, thereby effectively exploiting the correlation of temporal priors.
Our main contributions are summarized as follows:
We propose a fine-grained motion compression method for neural B-frame video coding, which involves an interactive dual-branch motion auto-encoder, per-branch adaptive quantization steps, and an interactive motion entropy model to reduce bi-directional motion coding costs.
We propose a selective temporal fusion method for neural B-frame video coding, which uses bi-directional fusion weights and hyperprior-based implicit alignment to enable discriminative utilization of bi-directional temporal contexts and priors in the contextual encoder-decoder and contextual entropy model.
Experimental results show that our codec outperforms state-of-the-art neural B-frame codecs and achieves comparable or even superior compression performance to the H.266/VVC reference software under random-access configuration.
We organize the remainder of this paper as follows. Section II reviews existing work on neural P-frame and B-frame video coding. Section III describes the details of our proposed methods. Section IV discusses experimental results and ablation studies, while Section V gives a conclusion for this paper.
II Related Work
II-A Neural P-Frame Video Coding
The majority of neural video codecs adopt the P-frame coding paradigm [4, 5, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 28, 24, 25, 26, 27, 29, 30, 6, 31, 35, 36, 37, 38, 39, 40, 41]. Lu et al. [4] proposed a pioneering work in this field—DVC [4], which implements key video coding components including motion estimation, motion compression, motion compensation, residual compression, and entropy models with neural networks and optimizes them in an end-to-end manner. Subsequently, Lin et al. [10] proposed to fuse the information of multiple previously compressed motion vectors to reduce the motion coding costs. Agustsson et al. [12] developed a scale-space flow, extending optical flow with a scale parameter to better model uncertainty, along with an enhanced warping operator robust to disocclusions and fast motion. Alternative to optical flow-based approaches, Hu et al. [11] proposed to represent motion with the offsets of deformable convolution and utilize deformable convolution to perform motion compensation. Based on this work, they further proposed to perform deformable convolution-based motion estimation and compensation in a coarse-to-fine manner [5].
The abovementioned schemes adopt a residual coding paradigm to utilize predicted temporal contexts. Beyond residual coding, Li et al. [27] proposed a conditional coding paradigm to better utilize temporal correlation, which fuses predicted temporal contexts into a contextual encoder-decoder and a contextual entropy model, enabling automatic learning of more effective methods for reducing temporal redundancy. Based on this scheme, Sheng et al. [31] proposed to replace reference frames with reference features to propagate temporal information and designed a temporal context mining module that performs flow-based warping on these reference features to learn multi-scale temporal contexts. This method leverages spatio-temporal non-uniformity to generate more accurate temporal predictions. Li et al. [35] further proposed a multi-granularity quantization mechanism to perform spatial-channel-wise quantization. This content-adaptive quantization mechanism not only helps achieve the smooth rate adjustment in a single model but also improves the compression performance by dynamic bit allocation. Through integration of these advanced techniques, recent DCVC-based learned P-frame codecs [38, 36, 6] have outperformed the reference software of traditional video coding standard H.266/VVC [3] under low-delay configurations by a large margin.
II-B Neural B-Frame Video Coding
The development of neural B-frame video coding [42, 43, 13, 44, 45, 46, 47, 48, 32, 49, 50, 51, 52, 33, 53, 54] is relatively slower compared with neural P-frame video coding. Based on the need for motion coding, existing neural B-frame coding schemes can be categorized into two types.
Schemes in the first category [42, 43, 13, 44] depend on video interpolation to perform bi-directional temporal prediction without the need for motion coding. Wu [42] et al. designed a contextual video interpolation network that uses bi-directional multi-scale contexts to obtain predicted frames. Following a similar idea, Cheng et al. [13] proposed to adaptively select the frame number in one interpolation loop according to video motion patterns. Xu [44] proposed to select interpolated temporal contexts adaptively using a residual-guided masking encoder. Alexandre et al. [43] introduced a two-layer neural B-frame codec that merges the low-resolution interpolated frames from the base layer and high-resolution interpolated frames from the enhancement layer to generate more accurate predicted frames. Currently, the best interpolation-based neural B-frame video codec [43] has outperformed x265—the industrial software of H.265/HEVC.
Schemes in the second category [45, 46, 47, 48, 32, 49, 50, 51, 52, 33, 53, 54] aim to enhance bi-directional temporal prediction by compressing bi-directional motion information. For example, Yang et al. [45] proposed to perform bi-directional motion vector estimation and compression to obtain bi-directional predicted frames. Subsequently, they employed a residual codec to compress the prediction residuals. Feng et al. [49] proposed to replace the bi-directional optical flows with a voxel flow and replace the bi-directional prediction merging process with the trilinear warping operation. Recently, some schemes [50, 32, 53, 33, 54] proposed to extend conditional coding [27] from P-frame to B-frame. For example, Sheng et al. [33] proposed DCVC-B that predicts bi-directional multi-scale temporal contexts using a bi-directional temporal context mining module. These contexts are then directly fused into a contextual encoder-decoder and a contextual entropy model to reduce temporal redundancy. Although DCVC-B [33] surpassed the reference software of H.265/HEVC standard with random-access configurations, it is difficult to achieve higher B-frame compression performance by simply applying P-frame coding methods to B-frame coding without fully considering the unique characteristics of bi-directional prediction.
III Methodology
The overview of our neural B-frame video codec is illustrated in Fig. 1. It adopts an architecture similar to [33]. To reduce bi-directional motion coding costs and achieve selective utilization of bi-directional temporal predictions, we propose a fine-grained motion compression method and a selective temporal fusion method.
III-A Fine-Grained Motion Compression
We employ SpyNet [55] to estimate bi-directional motion vectors (, ) and associated predictions (, ). Our fine-grained motion compression method is used to encode and decode their motion vector differences (, ) [33].
III-A1 Interactive Dual-Branch Motion Auto-Encoder
Our method incorporates an interactive dual-branch motion auto-encoder, illustrated in Fig. 2. Unlike the scheme that concatenated bi-directional motion vector differences (MVDs) for joint compression [33], our design processes forward and backward MVDs (, ) separately. This is motivated by the observation that bi-directional reference frames often exhibit asymmetric temporal correlations, necessitating distinct motion reconstruction qualities for each direction. The dual-branch motion encoder—comprising downsampling residual blocks and depth-wise convolution blocks—compresses MVDs into two latent representations (, ). These representations are quantized and entropy encoded into bitstreams. During decoding, the dual-branch motion decoder—comprising upsampling residual blocks and depth-wise convolution blocks—reconstructs the quantized latent representations (, ) into motion vectors (, ).
Unlike the scheme that employed a parameter-shared motion auto-encoder for separate bi-directional motion compression [34], our dual-branch motion auto-encoder designs a motion information interaction (MII) module to exploit the inherent geometric consistency between forward and backward motion vectors. As shown in Fig. 2, the motion information interaction module takes intermediate bi-directional motion features (, ) as inputs and outputs (, ). At the start of the MII module, two depth-wise convolution blocks are separately applied to (, ) to generate bi-directional query, key, and value features (), ().
Then we use an efficient attention mechanism [56] to exchange the bi-directional motion information, obtaining the initial cross-directional motion contexts ().
where is the softmax function. Subsequently, we obtain the refined motion contexts by screening () according to (, ).
where and are two depth-wise convolution blocks. Finally, and are transformed to more compact motion feature and conditioned on the cross-directional motion contexts .
where and are two depth-wise convolution blocks. With our proposed MII module, the information from dual motion branches can be exchanged, thereby exploiting the cross-directional motion dependencies and reducing the motion coding costs.
III-A2 Per-Branch Adaptive Quantization Steps
Adaptive quantization step techniques [35, 36, 38] have proven effective for precise rate-distortion control in neural video coding. While recent work [33] successfully extended this method from P-frame to B-frame coding, it employed identical quantization steps for bi-directional motion vectors. This uniform treatment fails to account for the distinct bitrate allocation needs and reconstruction quality requirements inherent to bi-directional motion vectors. To address this limitation, we propose to integrate independent learnable quantization steps into each branch of both the motion encoder and decoder, as presented in Fig. 2. One quantization step is composed of a global quantization step and a channel-wise quantization step. The global quantization step is a number related to the target bit rate. The global quantization step is a one-dimensional vector, and the length of the vector is the same as the number of channels of the quantized feature.
where refers to element-wise multiplication. This design enables fine-grained, direction-specific rate-distortion control for forward and backward motion vectors, optimizing compression efficiency for B-frame coding.
III-A3 Interactive Motion Entropy Model
Given the quantized bi-directional motion latent representations (, ), we propose an interactive motion entropy model to estimate their probability distributions for arithmetic coding. First, we derive the motion hyperprior and motion temporal prior through:
where and are motion hyper encoder and motion hyper decoder, respectively. is the motion temporal prior extractor [33]. Next, we partition and into four segments via spatial-channel splitting, as illustrated in Fig. 3. To effectively capture bi-directional motion dependencies, when compressing the current motion latent segment, we not only fuse the motion hyperprior, motion temporal prior, and previously compressed latent segments in the same direction, but also fuse the previously compressed latent segments in the other direction.
By interactively fusing bi-directional latent segments as directional priors, our entropy model achieves fine-grained motion dependency modeling and compression efficiency.
III-B Selective Temporal Fusion
Given the reconstructed bi-directional motion vectors, we employ the temporal context mining moulde [31] to extract extract multi-scale temporal contexts from reference features (), , yielding forward contexts and and backward contexts , as illustrated in Fig. 1. These contexts and previously compressed contextual latent representations () are fused into the contextual encoder-decoder and context entropy model to reduce temporal redundancy. Our selective temporal fusion method focuses on how to discriminative utilize these bi-directional temporal contexts and temporal priors.
III-B1 Bi-Directional Weighting-based Context Fusion
Our method first incorporates a bi-directional weighting fusion mechanism to build a discriminative contextual encoder-decoder. As illustrated in Fig. 4, unlike the schemes [33] that blindly fused bi-directional temporal contexts, our discriminative contextual encoder-decoder predicts bi-directional fusion weights to selective fuse the intermediate contextual feature with the bi-directional temporal contexts ().
where is a convolutional layer and refers to element-wise multiplication. Through the end-to-end optimization, the predicted fusion weights adaptively recalibrate the contributions of bi-directional temporal contexts based on their relevance to the target frame, thereby enabling selective temporal fusion that suppresses redundant or noisy temporal contexts.
III-B2 Implicit Alignment-based Prior Fusion
Previous schemes [33] tended to fuse the compressed contextual latent representations (, ) and small resolution of temporal contexts (, ) into the contextual entropy model as a temporal prior . However, these schemes ignore the temporal misalignment of latent representations and the predicted errors of temporal contexts, which fail to optimally exploit temporal correlations. Therefore, our method further incorporates a simple hyperprior-based implicit alignment for the contextual entropy model. Specifically, we first use two depth-wise convolution blocks to generate query, key, and value features () from hyperprior and temporal prior .
Considering the hyperprior is an estimation of the current contextual latent representation , we treat it as a surrogate to help align the temporal prior using an efficient attention mechanism [56].
Then, we output the refined temporal prior using the aligned feature.
where is a depth-wise convolution block. The aligned temporal prior is then fused with the hyperprior and spatial prior to achieve discriminative entropy modeling.
| MCL-JCV | UVG | HEVC Class B | HEVC Class C | HEVC Class D | HEVC Class E | Average | |
| HM-RA-GOP16 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| HM-LDB-GOP4 | 23.9 | 29.8 | 23.7 | 26.2 | 29.7 | 32.2 | 27.6 |
| VTM-RA-GOP32 | –38.0 | –35.9 | –39.1 | –33.7 | –30.4 | –37.9 | –35.8 |
| VTM-RA-GOP16 | –36.9 | –34.0 | –37.2 | –32.1 | –29.2 | –35.4 | –34.1 |
| VTM-LDB-GOP8 | –13.3 | –5.3 | –11.7 | –8.6 | –3.9 | –10.4 | –8.9 |
| DCVC-DC-GOP32 | –22.8 | –25.0 | –22.0 | –19.9 | –32.3 | –24.6 | –24.4 |
| DCVC-FM-GOP32 | –20.7 | –23.7 | –20.5 | –19.3 | –31.1 | –30.5 | –24.3 |
| B-CANF-GOP32 | 22.6 | 21.0 | 13.9 | 30.9 | 5.1 | 28.5 | 20.3 |
| DCVC-B-GOP32 | –16.8 | –18.0 | –26.8 | –15.8 | –36.7 | –42.9 | –26.1 |
| Ours-GOP32 | –27.5 | –27.4 | –34.7 | –28.4 | –43.5 | –49.0 | –35.1 |
IV Experiments
IV-A Experimental Setup
IV-A1 Training Datasets
For training, we employ the training split of the Vimeo-90k dataset [57], which comprises 64610 video clips at 448256 resolution. Since each clip contains only 7 consecutive frames, we utilize this dataset for training stages requiring sequence lengths of 3, 5, and 7 frames [33]. For longer sequence requirements (17 and 33 frames), we utilize the extended Vimeo collection introduced in [33], which offers 9000 higher-resolution (640360) sequences with 33 frames per clip. The detailed training strategy can be referred to [33].
IV-A2 Evaluation Datasets
The compression performance is evaluated on three standard benchmark datasets: the MCL-JCV dataset [58] containing 30 diverse 19201080 sequences, the UVG dataset [59] with 7 challenging 1080p large-motion sequences, and the HEVC common test sequences [60] comprising 16 videos (Classes B, C, D, E) at resolutions spanning from 416240 to 19201080. We compress 96 (2 complete GOPs and 1 incomplete GOP) and 97 (3 complete GOPs) consecutive frames of each test sequence to verify the generalization of our codec for complete GOP and incomplete GOP. The detailed GOP structures can be referred to [33].
IV-A3 Implementation Details
During training, the Lagrange multiplier in the loss function is configured as (85, 170, 380, 840) to measure the relative importance of distortion and bit rate. At inference time, our codec achieves different rate points by adjusting its learned quantization steps. The implementation is built in PyTorch with CUDA acceleration, trained using the AdamW optimizer [61] with default hyperparameter settings. The batch size is set to 8.
IV-A4 Evaluation Metrics
We compute Peak Signal-to-Noise Ratio (PSNR) for quality assessment, and use bits per pixel (bpp) to measure the bitrate cost. For comparative analysis between different codecs, we utilize the Bjøntegaard-Delta rate (BD-rate) metric, which calculates the average bitrate difference (%) relative to an anchor codec. Negative BD-rate values indicate bitrate savings, positive values denote increased bitrate requirements.
| MCL-JCV | UVG | HEVC Class B | HEVC Class C | HEVC Class D | HEVC Class E | Average | |
| HM-RA-GOP16 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| VTM-RA-GOP32 | –38.4 | –35.8 | –38.5 | –33.6 | –30.3 | –37.3 | –35.7 |
| VTM-RA-GOP16 | –35.8 | –32.4 | –35.2 | –31.2 | –28.3 | –33.5 | –32.7 |
| B-CANF-GOP32 | 18.9 | 14.0 | 7.0 | 24.3 | –0.7 | 10.2 | –12.3 |
| DCVC-B-GOP32 | –20.4 | –22.5 | –31.1 | –19.7 | –39.2 | –45.4 | –29.7 |
| Ours-GOP32 | –30.5 | –31.4 | –38.5 | –31.1 | –45.4 | –50.8 | –38.0 |
IV-B Quantitative Evaluation Results
IV-B1 Comparison with Existing Neural Video Codecs
To benchmark the compression performance of our proposed codec, we first conduct a comprehensive comparison with state-of-the-art neural video codecs, including both uni-directional inter-frame prediction-based P-frame codecs (DCVC-DC [35] and DCVC-FM [38]) and bi-directional prediction-based B-frame codecs (B-CANF [32] and DCVC-B), with all codecs configured using a 32-frame intra-period. When testing 96 frames, our rate-distortion curves presented in Fig. 6 demonstrate consistent superiority over competing neural video codecs. Quantitative results in Table I reveal that our codec achieves an average bitrate reduction of 35.1% compared to the HM-RA-GOP16 anchor, significantly outperforming both leading P-frame codecs (DCVC-DC: 24.4%, DCVC-FM: 24.3%) and B-frame codec DCVC-B (26.1%). Notably, on the MCL-JCV dataset, our codec provides more than 10% compression performance gain over DCVC-B. Similar superiority is achieved when testing 97 frames, as presented in Fig. 7 and Table II, establishing new state-of-the-art neural B-frame coding performance.
IV-B2 Comparison with Traditional Video Codecs
We further compare with traditional video codecs. Specifically, we include HM-16.20 [62] (H.265/HEVC reference software) tested under random-access (encoder_randomaccess_main_rext, HM-RA-GOP16) and low-delay (encoder_lowdelay_vtm, HM-LDB-GOP4) configurations, along with VTM-13.2 (H.266/VVC reference software) evaluated under random-access (encoder_randomaccess_vtm_gop16, VTM-RA-GOP16 and encoder_randomaccess_vtm, VTM-RA-GOP32) and low-delay (encoder_lowdelay_vtm, VTM-LDB-GOP8) configurations. All codecs are configured with an intra-period of 32. When compressing 96 frames, as reported in Fig. 6 and Table I, our codec achieves comparable performance to VTM, with an average BD-rate reduction of –34.5% versus –34.1% (VTM-RA-GOP16) and –35.8% (VTM-RA-GOP32). When testing 97 frames, our codec even achieves higher performance than VTM, with an average BD-rate reduction of –38.0% versus –32.7% (VTM-RA-GOP16) and –35.7% (VTM-RA-GOP32). These results demonstrate the strong potential of neural B-frame codecs.
| Codecs | MACs/pixel | Parameter | Enc Time | Dec Time |
| HM-RA | — | — | 92.24s | 0.29s |
| VTM-RA | —- | — | 1144.70s | 0.37s |
| DCVC-DC | 1397.90K | 18.45M | 0.82s | 0.64s |
| DCVC-FM | 1180.77K | 17.02M | 0.74s | 0.53s |
| B-CANF | 3081.11K | 23.66M | 1.49 s | 1.06s |
| DCVC-B | 3004.52K | 21.40M | 1.19s | 0.99s |
| Ours | 3382.25K | 26.07M | 1.47s | 1.10s |
IV-C Qualitative Comparison Results
We also present the qualitative comparison results of our codec and DCVC-B in Fig. 8. We can observe that the wooden floor textures of our decoded BasketballDrill_832x480_50 sequence maintain richer granular details. The bee’s wings of our decoded HoneyBee_1920x1080_120fps_420_8bit_YUV sequence are reconstructed with sharper transparency and vein patterns. The results indicate that our codec can achieve better visual enhancements while maintaining competitive bitrates.
IV-D Runtime and Computational Complexity Comparison
Table III presents a comprehensive comparison of computational complexity and encoding/decoding time across codecs for a 1080p video frame. Traditional codecs (HM, VTM) were evaluated on an Intel(R) Xeon(R) Gold 5118 CPU, while neural codecs were tested on an NVIDIA RTX 3090 GPU. Compared to DCVC-B, our method exhibits increased computational cost (3004.52K 3382.25K MACs/pixel) and model size (21.40M 26.07M parameters), with encoding/decoding time rising from 1.19s 1.47s and 0.99s 1.10s per frame, respectively. While these increases enable superior rate-distortion performance over DCVC-B, further complexity optimization is required. Reducing computational overhead remains a critical focus for our future work.
| Model | IDMAE | PBAQS | IMEM | BWF | HIA | BD-rate |
| ✗ | ✗ | ✗ | ✗ | ✗ | 0.0 | |
| ✓ | ✗ | ✗ | ✗ | ✗ | –2.6 | |
| ✓ | ✓ | ✗ | ✗ | ✗ | –4.9 | |
| ✓ | ✓ | ✓ | ✗ | ✗ | –6.3 | |
| ✓ | ✓ | ✓ | ✓ | ✗ | –9.1 | |
| ✓ | ✓ | ✓ | ✓ | ✓ | –10.4 |
IV-E Ablation Studies
IV-E1 Effectiveness of Fine-Grained Motion Compression Method
We first conduct an ablation study to validate the effectiveness of our proposed fine-grained motion compression method. To reduce training time, we only use 7-frame video sequences to train the models in the ablation studies, instead of fine-tuning them with the long-sequence dataset [33]. Our ablation study shows substantial performance gains through progressive integration of key components, as evidenced in Table IV. The interactive dual-branch motion auto-encoder (IDMAE) forms the foundation with a 2.6% BD-rate improvement, while the per-branch adaptive quantization steps (PBAQS) further deliver a 2.3% gain (from –2.6% to –4.9% BD-rate) by enabling fine-grained rate allocation tailored to each motion direction. The interactive motion entropy model (IMEM) further brings another 1.7% gain (from –4.9% to –6.3% BD-rate), which effectively captures the inherent correlations between bi-directional motion latent representations for enhanced compression efficiency. We compare the motion coding costs of the model with and without our proposed method in Fig. 9. We can observe that our method can effectively reduce the bi-directional motion compression bitrate.
IV-E2 Effectiveness of Selective Temporal Fusion Method
Similarly, we continue to conduct another ablation study to validate the effectiveness of our selective temporal fusion method by progressively integrating its key components into the baseline Model . As presented in Table IV, the introduction of bi-directional weighting fusion (BWF) first improves the compression performance by 2.8% BD-rate reduction (from –6.3% to –9.1% BD-rate), followed by an additional 1.3% gain (reaching –10.4% BD-rate) when incorporating the hyperprior-based implicit alignment (HIA).
We further visualize the predicted bi-directional temporal contexts along with their corresponding fusion weights in Fig. 10. Taking the frame from BasketballDrill_832x480_50 as an instance, we can observe that the forward temporal context exhibits relatively higher prediction quality than the backward temporal context. As a result, our proposed BWF assigns larger fusion weights to the forward temporal context to preserve its temporal information more effectively, while assigning smaller fusion weights to the backward temporal context to mitigate its negative impacts. In addition, we also visualize the residuals between the contextual latent representation of the frame of PartyScene_832x480_50 and its bi-directional temporal priors before () and after () being processed by our proposed HIA. As illustrated in Fig. 11, our HIA can effectively reduce the residual magnitudes, indicating a significant decrease in the local misalignment of temporal priors.
V Conclusion
We first propose a fine-grained motion compression method for neural B-frame video coding. By leveraging an interactive dual-branch motion auto-encoder, per-branch adaptive quantization steps, and an interactive motion entropy model, the bi-directional motion coding costs are effectively reduced. Furthermore, we propose a selective temporal fusion method. Using bi-directional fusion weights and hyperprior-based implicit alignment, our method enables discriminative utilization of bi-directional temporal contexts and temporal priors in the contextual encoder-decoder and contextual entropy model. Experimental results verify that our proposed codec outperforms state-of-the-art neural B-frame codecs and achieves comparable or even superior compression performance to the H.266/VVC reference software under random-access configuration. In the future, we target to achieve a better rate-distortion-complexity trade-off for neural B-frame coding.
VI Acknowledgment
We would like to express our gratitude to Zhirui Zuo for his assistance in dataset processing.